DIRECT OBSERVATION OF THE c^{3}Σ_{u}^{+} STATE OF C_{2}

DAMIAN L. KOKKIN, NEIL J. REILLY, JENNA A. JOESTER, \underline{MASAKAZU NAKAJIMA}, KLAAS NAUTA, SCOTT H. KABLE, and TIMOTHY W. SCHMIDT, School of Chemistry, University of Sydney, NSW 2006, Australia.

The d^{3}Π_{g} − c^{3}Σ_{u}^{+} system of dicarbon was observed in a discharged supersonic jet of C_{2}H_{2}.^{a} A spectral survey was carried out based on the result of high-level ab initio calculation.^{b} LIF excitation spectra of the band system were recorded by detecting fluorescence of the Swan system (d^{3}Π_{g} − a^{3}Π_{u}). Several molecular constants of the c^{3}Σ_{u}^{+} state were experimentally determined from the rotationally resolved spectra. Characterization of the c^{3}Σ_{u}^{+} state of C_{2} is of importance to astrophysical models of the species.

aD. L. Kokkin et al., J. Chem. Phys. \underline{125}, 231101 (2006).

bD. L. Kokkin et al., J. Chem. Phys. \underline{126}, 084302 (2007).